I need help on the following word problem.
A commuter bus takes 2 hours to get downtown; an express bus, averaging 25 mph faster, takes 45 minutes to cover the same route. What is the average speed for the commuter bus?

Answer by jorel1380(3719)

3/4(n+25)=2n
3n/4+75/4=2n
75/4=5n/4
5n=75
n=15
2n=30
The route is 30 miles long each way-
60 miles/2-3/4 hours=60/11/4=240/11=21.81 mph average..